OpenStudy (leann01xxx):

Somebody please help me. This Odysseyware program is really stressing me out. :'( 1. Find the sum of (5x3 + 3x2 - 5x + 4) and (8x3 -5x2 + 8x + 9). 2. Add: 2a3 + 5a - 7 and 5a2 - 6a + 12 3. Subtract: 11a2 + 42a - 6 from 14a 2 - 24a + 7. 4. (-6.7x 2 + 2.3x y + 5.2) - (-14.9x 2 - 3.5x y + 7.1)

OpenStudy (automaticloveletter):

did you try google

OpenStudy (leann01xxx):

Yes and I cant find any math website that can help me with them.

OpenStudy (anonymous):

1 Distribute sum groups 5x3(8x3−5x2+8x+9)+3x2(8x3−5x2+8x+9)−5x(8x3−5x2+8x+9)+4(8x3−5x2+8x+9) 2 Distribute 40x6−25x5+40x4+45x3+3x2(8x3−5x2+8x+9)−5x(8x3−5x2+8x+9)+4(8x3−5x2+8x+9) 3 Distribute 40x6−25x5+40x4+45x3+24x5−15x4+24x3+27x2−5x(8x3−5x2+8x+9)+4(8x3−5x2+8x+9) 4 Distribute 40x6−25x5+40x4+45x3+24x5−15x4+24x3+27x2−40x4+25x3−40x2−45x+4(8x3−5x2+8x+9) 5 Distribute 40x6−25x5+40x4+45x3+24x5−15x4+24x3+27x2−40x4+25x3−40x2−45x+32x3−20x2+32x+36 6 Gather like terms 40x6+(−25x5+24x5)+(40x4−15x4−40x4)+(45x3+24x3+25x3+32x3)+(27x2−40x2−20x2)+(−45x+32x)+36 7 Simplify 40x6−x5−15x4+126x3−33x2−13x+36

OpenStudy (automaticloveletter):

(5x^3 + 3x^2 - 5x + 4) + (8x^3 - 5x^2 + 8x + 9)=5x^3 + 3x^2 - 5x + 4 + 8x^3 - 5x^2 + 8x + 9=13x^3 - 2x^2 + 3x + 13

OpenStudy (leann01xxx):

The answer that it gives is too short for those answers. Or whatever that is.

OpenStudy (automaticloveletter):

options?

OpenStudy (p0sitr0n):

usually how you do this is by adding same powers of x together, i.e. \[ 5x^3 + 3x^2 - 5x + 4+8x^3 -5x^2 + 8x + 9=(5x^3+8x^3)+(3x^2-5x^2)+(-5x+8x)+(4+9)=13x^3-2x^2+3x+13\]

OpenStudy (p0sitr0n):

However, if you know how to do these operations, and simply want to save time (or check your answers), you should try http://www.wolframalpha.com/

OpenStudy (automaticloveletter):

that website is good ^

OpenStudy (leann01xxx):

This is what the answer box looks like

OpenStudy (leann01xxx):

?

Nnesha (nnesha):

\(\huge\color{green}{{\rm combine~\color{reD}{LIKE}~terms}}\)

Nnesha (nnesha):

sum = add so \[\huge\rm 5x^3 + 3x^2 - 5x + 4 +5a^2 - 6a + 12\] add coefficient of same variables (same degree) combine constant terms

OpenStudy (misty1212):

that is one ugly answer box! what kind of system is this?

OpenStudy (leann01xxx):

Its called Odysseyware

OpenStudy (misty1212):

homer would have loved it i am sure

OpenStudy (misty1212):

the last one looks the most annoying want to do that one fast?

OpenStudy (leann01xxx):

The answer boxes are all the same for those questions. And I have no idea what to do to get answers that are small enough to fit in the boxes.

OpenStudy (misty1212):

sorry can't help you there

OpenStudy (leann01xxx):

Im stuck on number 9 too. Factor completely, then place the answer in the proper location on the grid. Write the answer in descending powers of x. 6x 4 + 15x 3 y 2 + 3x 2 y 3

OpenStudy (misty1212):

they have you working on a machine, let the machine do it

Nnesha (nnesha):

how many terms you can write in that box ??

Nnesha (nnesha):

http://prntscr.com/7g5md9 ur answer is not right

OpenStudy (leann01xxx):

Each question box is different. Sometimes it goes from like 4 boxes to like two rows of boxes. @Nnesha And it doesn't do it itself. I have to figure out the answer and drag the number and symbols to the purple and place them in order. @misty1212

Nnesha (nnesha):

alright that's right it depends on the answer \[\huge\rm 5x^3 + 3x^2 - 5x + 4 +5a^2 - 6a + 12\] combine like terms what did you get ??

OpenStudy (anonymous):

http://www.mathpapa.com/algebra-calculator.html